# UniRGB-IR: A Unified Framework for RGB-Infrared Semantic Tasks via Adapter Tuning

### v2

@@ III-B Multi-modal Feature Pool
Figure 4: Structure of supplementary feature injector (SFI) module. A gated fusion unit (GFU) is utilized to dynamically fuse the current and the last injected features.
III-C Supplementary Feature Injector
As shown in Figure 4, we propose a supplementary feature injector (SFI) module to adaptively introduce the contextual multi-scale features without altering the ViT structure. Since the sequence lengths of contextual multi-scale features $F_{mfp}$ and the ViT features $F_{vit}^{i}$ are different, to address this, we employ sparse attention (eg. [42] and [43]) to dynamically sample supplementary features from each scale. To be specific, we utilize the ViT features $F_{vit}^{i}\in\mathbb{R}^{\frac{HW}{16^{2}}\times D}$ as the query, and the contextual multi-scale features $F_{mfp}\in\mathbb{R}^{(\frac{HW}{8^{2}}+\frac{HW}{16^{2}}+\frac{HW}{32^{2}})\times D}$ as the key and value:
$\tilde{F}_{sfi}^{i}=Attention(LN(F_{vit}^{i}),LN(F_{mfp})),$ (2)
where $Attention(\cdot)$ is the attention layer and $LN(\cdot)$ is LayerNorm [44] which is intended to reduce modality discrenpancy. Furthermore, we adopt progressive injection to introduce contextual multi-scale features, which can balance the foundation model features and the injected features $F_{sfi}^{i}$. Thus, a gated fusion unit is explored to predict the fusion weight $z$ to gate $F_{sfi}^{i-1}$ and $\tilde{F}_{sfi}^{i}$ for dynamic fusion. Specifically, we concatenate the two features $F_{sfi}^{i-1}$ and $\tilde{F}_{sfi}^{i}$ and feed it into linear layer to predict the weight $z$. Then, $z$ and $1-z$ are used to fuse $F_{sfi}^{i-1}$ and $\tilde{F}_{sfi}^{i}$ features respectively. The final output features $F_{sfi}^{i}$ of SFI module can be formulated as:
$F_{sfi}^{i}=\begin{cases}\tilde{F}_{sfi}^{i},&i=1\\ {GFU}(\tilde{F}_{sfi}^{i},F_{sfi}^{i-1}),&i=2\ldots N\end{cases}$ (3)
where $GFU(\cdot,\cdot)$ is the gated fusion unit shown in Figure 4.
@@ III-C Supplementary Feature Injector
TABLE I: Comparison results (mAP, in%) on the FLIR and LLVIP dataset. The best results are highlighted in red and the second-best are highlighted in blue. “–” indicates that the author did not provide the corresponding results.
@@
SSD [26] IR 24.6 57.0 18.0 53.5 90.2 57.9
RetinaNet [45] IR 31.5 66.1 25.3 55.1 94.8 57.6
Faster R-CNN [28] IR 37.6 75.8 31.6 54.5 94.6 57.6
Cascade R-CNN (ViT-B) [46] IR 41.9 78.6 37.3 60.4 94.9 67.5
DDQ-DETR [47] IR 37.1 73.9 32.2 58.6 93.9 64.6
SSD [26] RGB 18.8 46.3 13.1 39.8 82.6 31.8
RetinaNet [45] RGB 21.9 51.2 15.2 42.8 88.0 34.4
Faster R-CNN [28] RGB 27.7 62.2 21.2 45.1 87.0 41.2
Cascade R-CNN (ViT-B) [46] RGB 33.3 69.3 26.2 51.7 90.3 54.7
DDQ-DETR [47] RGB 30.9 64.9 24.5 46.7 86.1 45.8
GAFF [48] RGB+IR 37.4 74.6 31.3 55.8 94.0 60.2
ProbEn [49] RGB+IR 37.9 75.5 31.8 51.5 93.4 50.2
LGADet [50] RGB+IR - 74.5 - - - -
CSAA [51] RGB+IR 41.3 79.2 37.4 59.2 94.3 66.6
@@
TABLE II: Comparison results ($\text{MR}^{\text{-2}}$, in%) under ‘All-dataset’ settings of different pedestrian distances, occlusion levels, and light conditions (Day and Night) on the KAIST dataset. The best and second results are highlighted in red and blue respectively.
@@
ACF [52] 28.74 53.67 88.20 62.94 81.40 88.08 64.31 75.06 67.74
Halfway Fusion [53] 8.13 30.34 75.70 43.13 65.21 74.36 47.58 52.35 49.18
FusionRPN+BF [54] 0.04 30.87 88.86 47.45 56.10 72.20 52.33 51.09 51.70
IAF R-CNN [55] 0.96 25.54 77.84 40.17 48.40 69.76 42.46 47.70 44.23
IATDNN+IASS [56] 0.04 28.55 83.42 45.43 46.25 64.57 49.02 49.37 48.96
CIAN [25] 3.71 19.04 55.82 30.31 41.57 62.48 36.02 32.38 35.53
MSDS-R-CNN [57] 1.29 16.19 63.73 29.86 38.71 63.37 32.06 38.83 34.15
AR-CNN [27] 0.00 16.08 69.00 31.40 38.63 55.73 34.36 36.12 34.95
MBNet [58] 0.00 16.07 55.99 27.74 35.43 59.14 32.37 30.95 31.87
TSFADet [16] 0.00 15.99 50.71 25.63 37.29 65.67 31.76 27.44 30.74
CMPD [59] 0.00 12.99 51.22 24.04 33.88 59.37 28.30 30.56 28.98
CAGTDet [29] 0.00 14.00 49.40 24.48 33.20 59.35 28.79 27.73 28.96
C2Former [5] 0.00 13.71 48.14 23.91 32.84 57.81 28.48 26.67 28.39
@@
III-D Adapter Tuning Paradigm
To fully inherit the prior knowledge of the ViT pre-trained on large-scale datasets, we explore the adapter tuning paradigm instead of the full fine-tuning manner. For the dataset $D=\{(x_{j},{gt}_{j})\}_{j=1}^{M}$ of the downstream task, full fine-tuning process calculates the loss between the prediction and the ground truth, which can be formulated as:
$\mathcal{L}(D,\theta)=\sum_{j=1}^{M}\operatorname{loss}(F_{\theta}(x_{j}),{gt}_{j}),$ (4)
@@ III-D Adapter Tuning Paradigm
$\mathcal{L}(D,{\theta_{V}},{\theta_{A}})=\sum_{j=1}^{M}\operatorname{loss}(F_{\theta_{V},\theta_{A}}(x_{j}),{gt}_{j}),$ (6)
$\theta_{A}\leftarrow\underset{\theta_{A}}{\arg}\min\mathcal{L}(D,\theta_{V},\theta_{A}).$ (7)
@@
TABLE III: Results on the PST900 dataset. The best results are highlighted in red and the second-best are highlighted in blue.
“–” indicates that the author did not provide the corresponding results.
Methods Background Fire-Extinguisher Backpack Hand-Drill Survivor mAcc mIoU
Acc IoU Acc IoU Acc IoU Acc IoU Acc IoU
MFNet [18] 99.9 98.7 71.8 67.4 52.8 52.4 46.7 39.3 18.8 18.9 58.0 55.3
RTFNet [60] 99.9 99.0 78.6 54.8 62.5 60.8 76.7 61.0 65.2 62.5 76.6 67.6
CCNet [61] 99.6 98.7 88.1 73.8 76.0 73.0 54.1 51.0 49.5 33.5 73.46 66.0
PSTNet [62] - 98.9 - 70.1 - 69.2 - 53.6 - 50.0 - 68.4
ACNet [63] 99.8 99.3 84.9 60.0 85.6 83.2 53.6 51.5 69.1 65.2 78.7 71.8
FDCNet [64] 99.7 99.2 91.8 71.5 77.5 72.2 82.5 70.4 78.4 72.4 86.0 77.1
CGFNet [65] 99.7 99.3 96.3 71.7 87.6 82.0 70.5 59.7 86.3 77.4 88.1 78.0
EGFNet [66] 99.5 99.2 95.2 71.3 94.2 83.1 98.0 64.7 83.3 74.3 94.0 78.5
CCFFNet [30] 99.9 99.4 87.0 79.3 80.2 77.8 88.3 77.4 77.9 73.4 86.7 81.5
RSFNet [67] - 99.4 - 75.4 - 84.9 - 72.9 - 70.1 - 80.5
SGFNet [68] 99.8 99.4 89.4 75.6 90.4 85.4 94.0 76.7 82.7 76.7 91.2 82.8
IV Object Detection and Semantic Segmentation
To evaluate the effectiveness of our UniRGB-IR, we utilize the ViT-Base model (pre-trained on COCO [10] dataset) as the foundation model and utilize this framework to perform different RGB-IR semantic tasks. During training, we freeze the ViT-Base model and only optimize the MFP and SFI modules. We evaluate and compare our method with various competitive models, including CNN-based and Transformer-based models. Besides, our evaluation spans various tasks, including RGB-IR object detection on FLIR [69], LLVIP [70], and KAIST [52] datasets, RGB-IR semantic segmentation on MFNet [18] and PST9000 [62] datasets. Furthermore, ablation experiments on the designed modules and qualitative experiments are also conducted to verify that the UniRGB-IR framework can be leveraged as a unified framework to efficiently introduce IR features into the foundation model to achieve superior performance.
Figure 5: The Miss rate curves for all detectors are plotted against False Positives Per Image (FPPI) under the IoU threshold of 0.5 on the KAIST dataset.
IV-A RGB-IR Object Detection
Datasets. Our object detection experiments are based on the three paired RGB and IR object detection datasets. FLIR [69] is a paired visible and infrared object detection dataset, including daytime and night scenes, which has 4,129 aligned RGB-IR image pairs for training and 1,013 for testing. For LLVIP [70] dataset, it contains 15,488 aligned RGB-IR image pairs, of which 12,025 images are used for training and 3,463 images for testing. As for KAIST [52] dataset, it is a aligned multispectral pedestrian deteciton dataset, in which 8,963 and 2,252 image pairs are utilized for training and testing.
Metrics. For FLIR and LLVIP datasets, we employ mean Average Precision (mAP) to evaluate the detection performance. As for KAIST dataset, we use log-average miss rate MR over the false positive per image (FPPI) with the range of [$10^{-2}$, $10^{0}$] to evaluate the pedestrian detection performance.
Settings. All the experiments are conducted with NVIDIA GeForce RTX 3090 GPUs. We implement our framework on the MMDetection library and use the Cascade R-CNN as the basic framework to perform RGB-IR object detection. The detector is trained with an initial learning rate of 2$\times$10-4 for 48 epochs. The batch size is set to 16, and the AdamW optimizer is employed with a weight decay of 0.1. Horizontal flipping is also used for data augmentation.
@@ IV-A RGB-IR Object Detection
Results on FLIR and LLVIP datasets. We compare our method with five common mono-modality methods and four competitive multi-modality methods. As shown in Table I, it can be seen that most of the multi-modality detectors are even worse than mono-modality detectors (eg. Faster R-CNN in IR modality). Since the RGB feature interfere with infrared feature under limited illumination conditions, it has a negative impact on the fused features utilized for object detection tasks. However, our UniRGB-IR can effectively solve this problem through the SFI module, enabling the detector to achieve better classification and localization processes.
@@
Results on KAIST dataset. The quantitative results of the different methods on the KAIST dataset are shown in Table II. The experiments are conducted under ‘All-dataset’ settings [52]. We compare our UniRGB-IR with thirteen multi-modal object detection methods. Our model achieves the best performance on the ‘All’, ‘Day’, and ‘Night’ conditions and the other four of five subsets (‘Near’, ‘Far’, ‘None’, ‘Partial’ and ‘Heavy’), and rank second in the ‘Medium’ subset. Furthermore, our detector surpasses the previous best competitor C2Former by 3.18% on the ‘All’ condition, which indicates UniRGB-IR is robust to the complex scenes. Intuitively, we draw the log-average Miss Rate (MR) over the False Positive Per Image (FPPI) curve of these detectors as shown in Figure 5. All these results verify that our UniRGB-IR can exploit the potential representation from the RGB-based foundation model to achieve superior results.
TABLE IV: Results on the MFNet dataset. The best results are highlighted in red and the second-best are highlighted in blue.
“–” indicates that the author did not the corresponding results.
Methods Unlabeled Car Person Bike Curve Car Stop Guardrail Color Cone Bump mAcc mIoU
Acc IoU Acc IoU Acc IoU Acc IoU Acc IoU Acc IoU Acc IoU Acc IoU Acc IoU
MFNet [18] 98.7 96.9 77.2 65.9 67.0 58.9 53.9 42.9 36.2 29.9 12.5 9.9 0.1 0.0 30.3 25.2 30.0 27.7 45.1 39.7
RTFNet [60] 99.4 98.5 93.0 87.4 79.3 70.3 76.8 62.7 60.7 45.3 38.5 29.8 0.0 0.0 45.5 29.1 74.7 55.7 63.1 53.2
PSTNet [62] 98.8 97.4 89.4 77.0 74.8 67.2 65.8 50.8 42.5 35.2 34.6 25.2 0.0 0.0 32.7 29.9 61.1 55.0 55.5 48.6
FuseSeg [71] 99.0 97.6 93.1 87.9 81.4 71.7 78.5 64.6 68.4 44.8 29.1 22.7 63.7 6.4 55.8 46.9 66.4 47.9 70.6 54.5
ABMDRNet [72] 98.6 97.8 94.3 84.4 90.0 69.6 75.7 60.3 64.0 45.1 44.1 33.1 31.0 5.1 61.7 47.4 66.2 50.0 69.5 54.8
MFFENet [73] 99.3 97.8 91.4 87.1 82.6 74.4 76.7 61.3 58.7 45.6 44.9 30.6 60.0 5.2 64.4 57.0 72.7 40.5 72.3 55.5
EGFNet [66] 98.7 98.0 95.8 87.6 89.0 69.8 80.6 58.8 71.5 42.8 48.7 33.8 33.6 7.0 65.3 8.3 71.1 47.1 72.7 54.8
MTANet [74] 98.4 98.0 95.8 88.1 90.9 71.5 80.3 60.7 75.3 40.9 62.8 38.9 38.7 13.7 63.8 45.9 70.8 47.2 75.2 56.1
MLFNet [75] - 97.3 - 82.3 - 68.1 - 67.3 - 27.3 30.4 - 15.7 - 55.6 - 40.1 - 53.8
FEANet [76] 99.3 98.2 93.3 87.8 82.7 71.1 76.7 61.1 65.5 46.5 26.6 22.1 70.8 6.6 66.6 55.4 77.3 48.9 73.2 55.3
GMNet [77] 99.2 97.5 94.1 86.5 83.0 73.1 76.9 61.7 59.7 44.0 55.0 42.3 71.2 14.5 54.7 48.7 73.1 47.4 74.1 57.3
CCFFNet [30] 98.8 98.3 94.5 89.6 83.6 74.2 73.2 63.1 67.2 50.5 38.7 31.9 30.6 4.8 55.2 49.7 72.9 56.3 68.3 57.6
CCAFFMNet [78] 97.4 95.2 95.0 88.6 86.1 72.9 82.1 67.1 71.2 45.9 32.1 24.8 57.1 17.8 58.0 50.1 76.2 57.8 72.8 57.8
CMX  [31] - 98.3 - 89.4 - 74.8 - 64.7 - 47.3 - 30.1 - 8.1 - 52.4 - 59.4 - 58.2
FDCNet [64] 98.8 98.2 94.1 87.5 91.4 72.4 78.1 61.7 70.1 43.8 34.4 27.2 61.5 7.3 64.0 52.0 74.5 56.6 74.1 56.3
ECGFNet [79] 99.3 97.5 89.4 83.5 85.2 72.1 72.9 61.6 62.8 40.5 44.8 30.8 45.2 11.1 57.2 49.7 65.1 50.9 69.1 55.3
LASNet [80] 97.6 97.4 94.9 84.2 81.7 67.1 82.1 56.9 70.7 41.1 56.8 39.6 59.5 18.9 58.1 48.8 77.2 40.1 75.4 54.9
UniRGB-IR (Ours) 98.3 97.2 94.0 83.7 88.7 64.9 88.0 69.8 53.3 36.8 58.5 41.0 69.7 36.7 77.6 56.2 55.2 47.3 75.7 59.3

### v4

@@ 3.1. Overall Architecture
Figure 4. Structure of supplementary feature injector (SFI) module. A gating network is utilized to dynamically fuse the current and the last injected features.
@@
Table 1. RGB-IR object detection results (mAP, in%) on the FLIR and LLVIP dataset. The best results are highlighted in red and the second-best are highlighted in blue. “–” indicates that the authors do not provide the corresponding results.
@@
SSD (Liu et al., 2016a) IR 24.6 57.0 18.0 53.5 90.2 57.9
RetinaNet (Lin et al., 2017) IR 31.5 66.1 25.3 55.1 94.8 57.6
Faster R-CNN (Ren et al., 2015) IR 37.6 75.8 31.6 54.5 94.6 57.6
Cascade R-CNN (ViT-B) (Cai and Vasconcelos, 2018) IR 41.9 78.6 37.3 60.4 94.9 67.5
DDQ-DETR (Zhang et al., 2023b) IR 37.1 73.9 32.2 58.6 93.9 64.6
SSD (Liu et al., 2016a) RGB 18.8 46.3 13.1 39.8 82.6 31.8
RetinaNet (Lin et al., 2017) RGB 21.9 51.2 15.2 42.8 88.0 34.4
Faster R-CNN (Ren et al., 2015) RGB 27.7 62.2 21.2 45.1 87.0 41.2
Cascade R-CNN (ViT-B) (Cai and Vasconcelos, 2018) RGB 33.3 69.3 26.2 51.7 90.3 54.7
DDQ-DETR (Zhang et al., 2023b) RGB 30.9 64.9 24.5 46.7 86.1 45.8
GAFF (Zhang et al., 2021a) RGB+IR 37.4 74.6 31.3 55.8 94.0 60.2
ProbEn (Chen et al., 2022b) RGB+IR 37.9 75.5 31.8 51.5 93.4 50.2
CSAA (Cao et al., 2023) RGB+IR 41.3 79.2 37.4 59.2 94.3 66.6
ICAFusion (Shen et al., 2024) RGB+IR 41.4 79.2 36.9 - - -
RSDet (Zhao et al., 2024b) RGB+IR 43.8 83.9 40.1 61.3 95.8 70.4
@@
3.2. Multi-modal Feature Pool
To complement the rich feature representations for RGB-IR semantic tasks, we introduce a multi-modal feature pool (MFP) module, including multiple perception and feature pyramid. The former can extract the contextual features with different convolution kernels, which achieve the long-distance modeling capability of CNNs. Different from existing works (He et al., 2019; Wu et al., 2022b) that increase the width or depth of the model, we efficiently achieve multi-receptive field perception in the channel dimension. As for feature pyramid, it can obtain multi-scale features to enhance the small object features. Therefore, these two operations are connected in series, enabling the MFP module to efficiently provide rich RGB-IR feature representations for various visible infrared semantic tasks, as shown in Figure 3.
Specifically, for the input RGB ($H\times W\times 3$) and IR ($H\times W$) images, we first employ a stem block borrowed from ResNet (He et al., 2016) to extract two modality features $F^{rgb}_{1}$ and $F^{ir}_{1}\in\mathbb{R}^{H/4\times W/4\times C}$. Then, these two features are split into four equal parts by utilizing channel splitting. To extract multi-receptive field perception, each part is subjected to convolution operations with different kernel sizes ($\times 3, 3×3$$\times 3$, $5\times 5$ and $7\times 7$). Then, we fuse each processed feature from two modalities by using SE attention (Hu et al., 2018) (shown in Figure 3). Therefore, we concatenate each fused part to obtain RGB-IR contextual features $F_{fus}$, which can be formulated as:
(1) $F_{fus}=\varGamma_{k=1}^{4}(Fus(W_{k}^{rgb}*f_{k}^{rgb},W_{k}^{ir}*f_{k}^{ir})),$
where $F_{fus}\in\mathbb{R}^{H/4\times W/4\times C}$, $f_{k}^{rgb}$ and $f_{k}^{ir}$ are the $k$-th part of $F_{1}^{rgb}$ and $F_{1}^{ir}$ features respectively, $W_{k}$ is the convolution with $k$-th kernel size, $\varGamma$ is the concatenation operation, $Fus(\cdot,\cdot)$ denotes the fusion module as shown in Figure 3. For the feature pyramid, a stack of three $3\times 3$ convolutions with $stride=2$ is applied to downsample the size of the feature maps. Then, features of each scale are fed into a $1\times 1$ convolution to project the feature maps to $D$ dimensions. Therefore, we can obtain a set of multi-scale features $F_{2}$, $F_{3}$, $F_{4}$} with 1/8, 1/16, and 1/32 of the original image resolution, respectively. Finally, we flatten and concatenate these features into feature tokens $F_{mfp}\in\mathbb{R}^{(\frac{HW}{8^{2}}+\frac{HW}{16^{2}}+\frac{HW}{32^{2}})\times D}$, which will be used as supplementary features for the ViT foundation model.
3.3. Supplementary Feature Injector
To adaptively introduce the contextual multi-scale features without altering the ViT structure, we propose a supplementary feature injector (SFI) module as shown in Figure 4. Since the sequence lengths of contextual multi-scale features $F_{mfp}$ and ViT features $F_{vit}^{i}$ are different, to address this, we employ sparse attention (eg. Pale Attention (Wu et al., 2022c) and Deformable Attention (Zhu et al., 2020)) to dynamically sample supplementary features from each scale. Specifically, we utilize the ViT features $F_{vit}^{i}\in\mathbb{R}^{\frac{HW}{16^{2}}\times D}$ as the query, and the contextual multi-scale features $F_{mfp}\in\mathbb{R}^{(\frac{HW}{8^{2}}+\frac{HW}{16^{2}}+\frac{HW}{32^{2}})\times D}$ as the key and value, which can be represented as:
(2) $\tilde{F}_{sfi}^{i}=Attention(LN(F_{vit}^{i}),LN(F_{mfp})),$
where $Attention(\cdot)$ is the sparse attention and $LN(\cdot)$ is LayerNorm (Ba et al., 2016), which aims to reduce modality differences during training. Furthermore, we adopt progressive injection to introduce contextual multi-scale features, which can balance the foundation model features and the injected features $F_{sfi}^{i}$. Thus, a gating network is explored to predict the fusion weight $z$ to gate $F_{sfi}^{i-1}$ and $\tilde{F}_{sfi}^{i}$ for dynamic fusion. Specifically, we concatenate the two features $F_{sfi}^{i-1}$ and $\tilde{F}_{sfi}^{i}$ and feed it into linear layer to predict the weight $z$. Then, $z$ and 1z1-z are used to fuse $F_{sfi}^{i-1}$ and $\tilde{F}_{sfi}^{i}$ features respectively. The final output features $F_{sfi}^{i}$ of SFI module can be formulated as:
(3) $F_{sfi}^{i}=\begin{cases}\tilde{F}_{sfi}^{i}\ ,&i=1\\ (1-z)*\tilde{F}_{sfi}^{i}+z*F_{sfi}^{i-1}\ .&i=2\ldots N\end{cases}$
3.4. Adapter Tuning Paradigm
To fully inherit the prior knowledge of the ViT pre-trained on large-scale datasets, we explore the adapter tuning paradigm instead of the full fine-tuning manner. For the dataset $D=\{(x_{j},{gt}_{j})\}_{j=1}^{M}$ of the different semantic tasks, full fine-tuning process calculates the loss between the prediction and the ground truth, which can be formulated as:
(4) $\mathcal{L}(D,\theta)=\sum_{j=1}^{M}\operatorname{loss}(F_{\theta}(x_{j}),{gt}_{j}),$
(5) $\theta\leftarrow\underset{\theta}{\arg\min}\mathcal{L}(D,\theta).$
(6) $\mathcal{L}(D,{\theta_{V}},{\theta_{A}})=\sum_{j=1}^{M}\operatorname{loss}(F_{\theta_{V},\theta_{A}}(x_{j}),{gt}_{j}),$
(7) $\theta_{A}\leftarrow\underset{\theta_{A}}{\arg}\min\mathcal{L}(D,\theta_{V},\theta_{A}).$
Table 2. RGB-IR pedestrian detection results ($\text{MR}^{\text{-2}}$, in%) under ‘All-dataset’ settings of different pedestrian distances, occlusion levels, and light conditions (Day and Night) on the KAIST dataset. The best and second results are highlighted in red and blue.
@@ 3.4. Adapter Tuning Paradigm
ACF (Hwang et al., 2015) 28.74 53.67 88.20 62.94 81.40 88.08 64.31 75.06 67.74
Halfway Fusion (Liu et al., 2016b) 8.13 30.34 75.70 43.13 65.21 74.36 47.58 52.35 49.18
FusionRPN+BF (Konig et al., 2017) 0.04 30.87 88.86 47.45 56.10 72.20 52.33 51.09 51.70
IAF R-CNN (Li et al., 2019) 0.96 25.54 77.84 40.17 48.40 69.76 42.46 47.70 44.23
IATDNN+IASS (Guan et al., 2019) 0.04 28.55 83.42 45.43 46.25 64.57 49.02 49.37 48.96
CIAN (Zhang et al., 2019b) 3.71 19.04 55.82 30.31 41.57 62.48 36.02 32.38 35.53
MSDS-R-CNN (Li et al., 2018) 1.29 16.19 63.73 29.86 38.71 63.37 32.06 38.83 34.15
AR-CNN (Zhang et al., 2019c) 0.00 16.08 69.00 31.40 38.63 55.73 34.36 36.12 34.95
MBNet (Zhou et al., 2020) 0.00 16.07 55.99 27.74 35.43 59.14 32.37 30.95 31.87
TSFADet (Yuan et al., 2022) 0.00 15.99 50.71 25.63 37.29 65.67 31.76 27.44 30.74
CMPD (Li et al., 2022b) 0.00 12.99 51.22 24.04 33.88 59.37 28.30 30.56 28.98
CAGTDet (Yuan et al., 2024) 0.00 14.00 49.40 24.48 33.20 59.35 28.79 27.73 28.96
C2Former (Yuan and Wei, 2024) 0.00 13.71 48.14 23.91 32.84 57.81 28.48 26.67 28.39
@@
Table 3. RGB-IR semantic segmentation on the PST900 dataset. The best results are highlighted in red and the second-best are highlighted in blue. “–” indicates that the authors do not provide the corresponding results.
Methods Background Fire-Exting. Backpack Hand-Drill Survivor mAcc mIoU
Acc IoU Acc IoU Acc IoU Acc IoU Acc IoU
MFNet (Ha et al., 2017) 99.9 98.7 71.8 67.4 52.8 52.4 46.7 39.3 18.8 18.9 58.0 55.3
RTFNet (Sun et al., 2019) 99.9 99.0 78.6 54.8 62.5 60.8 76.7 61.0 65.2 62.5 76.6 67.6
CCNet (Huang et al., 2019) 99.6 98.7 88.1 73.8 76.0 73.0 54.1 51.0 49.5 33.5 73.46 66.0
ACNet (Hu et al., 2019) 99.8 99.3 84.9 60.0 85.6 83.2 53.6 51.5 69.1 65.2 78.7 71.8
PSTNet (Shivakumar et al., 2020) - 98.9 - 70.1 - 69.2 - 53.6 - 50.0 - 68.4
FDCNet (Zhao and Zhang, 2022) 99.7 99.2 91.8 71.5 77.5 72.2 82.5 70.4 78.4 72.4 86.0 77.1
CCFFNet (Wu et al., 2022a) 99.9 99.4 87.0 79.3 80.2 77.8 88.3 77.4 77.9 73.4 86.7 81.5
RSFNet (Li et al., 2023a) - 99.4 - 75.4 - 84.9 - 72.9 - 70.1 - 80.5
MMDRNet (Liang et al., 2023) 99.5 98.9 77.9 52.4 79.1 71.1 77.1 40.6 69.8 62.3 81.3 68.7
SGFNet (Wang et al., 2023) 99.8 99.4 89.4 75.6 90.4 85.4 94.0 76.7 82.7 76.7 91.2 82.8
MDRNet (Zhao et al., 2024a) 99.3 99.1 90.2 63.0 93.1 76.3 86.6 63.5 85.6 71.3 91.0 74.6
KDSNet (Zhou et al., 2024) 99.8 99.4 92.0 82.3 83.2 81.2 86.1 71.3 72.6 70.5 86.8 80.9
4. Experiments
To evaluate the effectiveness of our UniRGB-IR, we utilize the ViT-Base model (pre-trained on COCO (Lin et al., 2014) dataset) as the foundation model and utilize this framework to perform RGB-IR semantic tasks. During training, we freeze the ViT-Base model and only optimize the MFP and SFI modules. We evaluate and compare our method with various competitive models, including CNN-based and Transformer-based models. Besides, our evaluation spans various tasks, including RGB-IR object detection on FLIR (Zhang et al., 2020), LLVIP (Jia et al., 2021), and KAIST (Hwang et al., 2015) datasets, RGB-IR semantic segmentation on PST900 (Shivakumar et al., 2020) and MFNet (Ha et al., 2017) (see supplementary materials) datasets and RGB-IR salient object detection on VT821 (Wang et al., 2018), VT1000 (Tu et al., 2019b) and VT5000 (Tu et al., 2022). Furthermore, ablation experiments on the designed modules and qualitative experiments are also conducted to verify that the UniRGB-IR framework can be leveraged as a unified framework to efficiently introduce RGB-IR features into the foundation model to achieve superior performance.
4.1. RGB-IR Object Detection
Datasets. Our object detection experiments are based on the three paired RGB and IR object detection datasets. FLIR (Zhang et al., 2020) is a paired visible and infrared object detection dataset, including daytime and night scenes, which has 4,129 aligned RGB-IR image pairs for training and 1,013 for testing. For LLVIP (Jia et al., 2021) dataset, it contains 15,488 aligned RGB-IR image pairs, of which 12,025 images are used for training and 3,463 images for testing. As for KAIST (Hwang et al., 2015) dataset, it is a aligned multispectral pedestrian deteciton dataset, in which 8,963 and 2,252 image pairs are utilized for training and testing.
Metrics. For FLIR and LLVIP datasets, we employ mean Average Precision (mAP) to evaluate the detection performance. As for KAIST dataset, we use log-average miss rate MR over the false positive per image (FPPI) with the range of $^{-2}, 1001$$^{0}$] to evaluate the pedestrian detection performance.
@@ 4.1. RGB-IR Object Detection
Table 4. RGB-IR salient object detection on VT821, VT1000 and VT5000 datasets. * represents RGB-D SOD transformed into RGB-T SOD. The best results are highlighted in red and the second-best are highlighted in blue.
Model VT821 VT1000 VT5000
$\emph{S}\uparrow$ $\emph{adpE}\uparrow$ $\emph{adpF}\uparrow$ $\emph{MAE}\downarrow$ $\emph{S}\uparrow$ $\emph{adpE}\uparrow$ $\emph{adpF}\uparrow$ $\emph{MAE}\downarrow$ $\emph{S}\uparrow$ $\emph{adpE}\uparrow$ $\emph{adpF}\uparrow$ $\emph{MAE}\downarrow$
MMCI* (Chen et al., 2019) 0.763 0.784 0.618 0.087 0.886 0.892 0.803 0.039 0.827 0.859 0.714 0.055
TANet* (Chen and Li, 2019) 0.818 0.852 0.717 0.052 0.902 0.912 0.838 0.030 0.847 0.883 0.754 0.047
S2MA* (Liu et al., 2020) 0.811 0.813 0.709 0.098 0.918 0.912 0.848 0.029 0.853 0.864 0.743 0.053
JLDCF* (Fu et al., 2021) 0.839 0.830 0.726 0.076 0.912 0.899 0.829 0.030 0.861 0.860 0.739 0.050
MTMR (Wang et al., 2018) 0.725 0.815 0.662 0.109 0.706 0.836 0.715 0.119 0.680 0.795 0.595 0.114
M3S-NIR (Tu et al., 2019a) 0.723 0.859 0.734 0.140 0.726 0.827 0.717 0.145 0.652 0.780 0.575 0.168
SGDL (Tu et al., 2019b) 0.765 0.847 0.731 0.085 0.787 0.856 0.764 0.090 0.750 0.824 0.672 0.089
FMSF (Zhang et al., 2019a) 0.760 0.796 0.640 0.080 0.873 0.899 0.823 0.037 0.814 0.864 0.734 0.055
MIDD (Tu et al., 2021) 0.871 0.895 0.803 0.063 0.915 0.933 0.880 0.027 0.868 0.896 0.799 0.043
ADF (Tu et al., 2022) 0.810 0.842 0.717 0.077 0.910 0.921 0.847 0.034 0.864 0.891 0.778 0.048
LSNet (Zhou et al., 2023b) 0.877 0.911 0.827 0.070 0.924 0.936 0.887 0.022 0.876 0.916 0.827 0.036
UniTR (Guo et al., 2024) 0.873 0.892 0.827 0.033 0.929 0.941 0.885 0.019 0.883 0.926 0.839 0.032
@@
Results on FLIR and LLVIP datasets. We compare our method with five common mono-modality methods and four competitive multi-modality methods. As shown in Table 1, it can be seen that most of the multi-modality detectors are even worse than mono-modality detectors (eg. Cascade R-CNN in IR modality). Since RGB features interfere with infrared feature under limited illumination conditions, it has a negative impact on the fused features utilized for object detection tasks. However, our UniRGB-IR can effectively solve this problem through the SFI module, enabling the detector to achieve better classification and localization processes.
@@
Results on KAIST dataset. The quantitative results of the different methods on the KAIST dataset are shown in Table 2. The experiments are conducted under ‘All-dataset’ settings (Hwang et al., 2015). We compare our UniRGB-IR with thirteen multi-modal object detection methods. Our model achieves the best performance on the ‘All’, ‘Day’, and ‘Night’ conditions and the other four of five subsets (‘Near’, ‘Far’, ‘None’, ‘Partial’ and ‘Heavy’), and rank second in the ‘Medium’ subset. Furthermore, our detector surpasses the previous best competitor C2Former by 3.18% on the ‘All’ condition, which indicates UniRGB-IR is robust to the complex scenes.
4.2. RGB-IR Semantic Segmentation
Datasets. Our semantic segmentation experiments are performed on the public RGB-IR semantic segmentation dataset PST900 (Shivakumar et al., 2020). The PST900 dataset is divided into 597 pairs for training and 288 pairs for testing, containing five categories (background, fire extinguisher, backpack, hand drill, and survivor). The dataset is divided into three parts: training, validation, and testing in a ratio of 2:1:1.